OpenStudy (anonymous):

HELP!!!!!!! Dx If sin theta =2/3 , which of the following are possible? 1. cos theta= -sqrt5/3 and tan theta=2/ sqrt5 2. sec theta= 3/sqrt5 and tan theta= 2/sqrt5 3. sec theta= -3/2 and tan theta= 2/sqrt5 4. cos theta= sqrt5/3 and tan theta= 2/sqrt5

OpenStudy (lgbasallote):

draw a triangle...sin theta = 2/3 = opposite/hypotenuse  that means the third side is \(\sqrt {3^2 - 2^2} = \sqrt{ 9 -4} = \sqrt 5\) so your adjacent is \(\sqrt 5\) opposite = 2 hypotenuse = 3 i think from that data you can verify which is correct..

OpenStudy (anonymous):

So it's 2 and 4 right?

Directrix (directrix):

@JCortez93 --> That is correct.